# Defining Neural Network Architecture through Polytope Structures of Dataset

### v1

@@ C.2 A polytope-basis cover derived from a two-layer ReLU network
Now, we quantize these functions. With the similar idea of Lebesgue integration, we can approximate the convex function $\mathcal{N}_{+}({\bm{x}})$ by a simple function. Here, the simple function means a linear combination of indicator functions, importantly considered in mathematical field like Lebesgue theory (Rudin et al., 1976). Let $M:=\max_{{\bm{x}}\in{{\mathcal{D}}}}\mathcal{N}_{+}({\bm{x}})$. Then, for a given $\varepsilon>0$, we can approximate $\mathcal{N}_{+}({\bm{x}})$ by
@@
where $\mathcal{C}_{Q}$ is collection of polytopes defined by $C:=\{{\bm{x}}~{}|~{}\mathcal{N}_{+}({\bm{x}})<M-l\varepsilon\}$ for $l=0,1,\cdots$. This decomposition can be understood as quantization of $\mathcal{N}_{+}({\bm{x}})$ by slices with height $\varepsilon$. Moreover, the above approximation would be accurate as $\varepsilon\rightarrow 0$. Similarly, $\mathcal{N}_{-}({\bm{x}})$ can be approximated in the similar manner. The main idea to obtain a polytope-basis cover $\mathcal{C}$ from $\mathcal{N}$ is selecting sufficiently many $\varepsilon$’s to quantize $\mathcal{N}_{+}$ and $\mathcal{N}_{-}$.
More precisely, we construct a polytope-basis cover from the values of $\mathcal{N}$. If $\mathcal{C}$ is not a polytope-basis cover yet, we select an incorrectly classified data point $\hat{{\bm{x}}}$ with the smallest confidence value, i.e., $\hat{\bm{x}}:=\operatorname*{arg\,min}_{{\bm{x}}_{i}\in{{\mathcal{D}}}}|\mathcal{N}({\bm{x}}_{i})|$. Subsequently, we choose an intermediate value $c$ between $\mathcal{N}_{+}(\hat{{\bm{x}}})$ and $\mathcal{N}_{-}(\hat{{\bm{x}}})$ and add two polytopes $C_{+}:={\mathcal{N}_{+}<c}$ and $C_{-}:={\mathcal{N}_{-}>-c}$ to the cover $\mathcal{C}$. Then, the value
@@
        $\mathcal{C}\leftarrow\mathcal{C}_{P}\cup\mathcal{C}_{Q}$
@@
The result of Algorithm 3 on a network with architecture $2\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}50\rightarrow 1$, which is trained on the swiss roll dataset, is illustrated in Figure 8. Specifically, the algorithm generates a polytope-basis cover of the dataset consists of 68 polytopes ($\mathcal{C}_{P}$ and $\mathcal{C}_{Q}$ consists of 34 polytopes). 14 polytopes in $\mathcal{C}$ is illustrated in 2nd column to 8th column in Figure 8. Proposition C.2 guarantees that Algorithm 3 must terminate in finite time, and produces a polytope-basis cover of ${{\mathcal{D}}}$ which has the same accuracy with the $\mathcal{N}$. Therefore, training a two-layer ReLU network to 100% accuracy on the dataset ${{\mathcal{D}}}$, Algorithm 3 allows to derive a polytope-basis cover of the given dataset.
However, there are two drawbacks in this algorithm : 1. generally it induces many polytopes in practice, and 2. the number of faces of each polytope is unknown. For detail comparison with other algorithms, see Section C.4.
C.3 An efficient algorithm to find an optimal polytope-basis cover
In the preceding subsections, we introduced two algorithms in Sections C.1 and C.2 that extract a polytope-basis cover from trained two-layer or three-layer ReLU networks. However, the results obtained from both algorithms, as illustrated in Figure 8 and 2, still exhibit multiple polytopes on the training dataset. As demonstrated for the swiss roll dataset in Figure 1, we have previously shown the existence of a polytope-basis cover comprising only four polytopes (refer to Figure 1 and 2).
@@ C.3 An efficient algorithm to find an optimal polytope-basis cover
Before we demonstrate the algorithm, we provide modified network and loss functions first. Specifically, we consider the following two types of two-layer ReLU networks. Here, $\lambda_{bias}>0$ is a hyperparameter that enhances the convergences.
@@
If $\lambda_{bias}$ has large value, then $\mathcal{T}_{+}$ has large output value at initialization, and gradient descent optimization heavily affected from data with 00 labels. The similar situation happens for $\mathcal{T}_{-}$, and it helps to find a single polytope that contains whole class data. Practically, it is enough to use $\lambda_{bias}=5$ to find such polytopes.
@@
where $\bm{\lambda}=(\lambda_{0},\lambda_{1})$ is the hyperparameter proposed to reinforce to cover a whole data class with specific label. By using a large value of $\lambda_{1}$, $\mathcal{T}$ is trained to cover whole data points that have label $y_{i}=1$. After successfully configured the first polytope $C_{1}$, we train the second network $\mathcal{T}_{2}$ to distinguish data points of another data class, $y_{i}=0$, in the obtained polytope $C_{1}$. Repeating this process alternatively, Algorithm 4 generally provides a polytope-basis cover with a small number of polytopes.
Now, we provide a detailed illustration of the algorithm’s functionality with the example displayed in Figure 9. Here we use $\lambda_{bias}=5$, and $\bm{\lambda}=(1,10)$. First, the algorithm trains $\mathcal{T}_{1}$ on the entire dataset ${{\mathcal{D}}}$ using (12). After fine-tuning through Algorithm 1, we obtain the first polytope $C_{1}$ displayed in Figure 9(b). Note that all orange data points $({{\mathcal{D}}}_{1})$ are contained in $C_{1}$. Next, the second network $\mathcal{T}_{2}$ is trained on $({{\mathcal{D}}}_{0}\cap C_{1})\cup{{\mathcal{D}}}_{1}$ using (12) with $\bm{\lambda}=(1,10)$. $\mathcal{T}_{2}$ is aimed to cover all blue data points $({{\mathcal{D}}}_{0})$ within $C_{1}$. After training and fine-tuning, we obtain the second polytope $C_{2}$ displayed in Figure 9(c). Similarly, we can find the third polytope $C_{3}$ by training another network $\mathcal{T}_{3}$ on ${{\mathcal{D}}}_{0}\cup({{\mathcal{D}}}_{1}\cap C_{2})$, displayed in Figure 9(d) and so on. In the example in Figure 9, a total of four polytopes are obtained ((b) to (e)). Lastly, by Theorem 3.4, we can construct a three-layer ReLU network with architecture $2\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}23\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}4\rightarrow 1$ that can completely classify this swiss roll dataset, based on the obtained polytope-basis cover. The decision boundary of this network is illustrated in Figure 9(f).
Algorithm 4 The efficient algorithm for finding a polytope-basis cover
@@
The key advantage of this algorithm lies in its ability to generate a small number of convex polytopes, in contrast to other algorithms we proposed.As demonstrated in the figure above, it suggests $2\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}23\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}4\rightarrow 1$ as a feasible architecture of the given swiss roll dataset, which apears to be close to optimal. However, due to its iterative nature of training multiple two-layer networks until achieving a complete polytope-basis cover, it typically requires a longer computation time. A detailed comparison with other algorithms is provided in the subsequent section.
Figure 9: The result of Algorithm 4. For the given dataset (a), the algorithm determines the first polytope $C_{1}$ which contains whole orange class (b). In the next step, it obtains the second polytope $C_{2}$ which contains the whole blue class inside $C_{1}$. Other polytopes are similarly obtained and described in (c) to (e). Totally, the algorithm produces a polytope-basis cover consisting of four polytopes. Theorem 3.4 shows that $2\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}17\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}4\rightarrow 1$ is a feasible architecture on this dataset, and the decision boundary of the induced network induced is drawn in (f).
C.4 Comparison of proposed algorithms
Finally, we compare the proposed algorithms: Algorithm 2, 3, and 4. First, we provide theoretical results for the proposed algorithms. The proof involves demonstrating how a polytope-basis cover can be explicitly constructed from the results of the algorithms. The detailed proof is available in Appendix E.6.
@@ C.4 Comparison of proposed algorithms
    Algorithm 2 must terminate in finite time, and it produces a polytope-basis cover that has the same accuracy with the compressed network $\mathcal{N}$.
@@
    Algorithm 3 must terminate in finite time, and it produces a polytope-basis cover that has the same accuracy with the given network $\mathcal{N}$.
@@
# of obtained polytopes at most $J$ large small
@@
Accuracy of the polytope cover unknown same with the given network 100% (if it terminates)
Table 3: Comparison of proposed algorithms. These algorithms all generate a polytope-basis cover of the given training dataset, but each of them has its own advatage and disadvantage.
@@
  • Input networks. Algorithm 2 operates on a pre-trained three-layer network outlined in (6). Algorithm 3 necessitates a fully-trained two-layer ReLU network specified by (1). However, Algorithm 4 does not necessitate any pre-existing networks as inputs and generates polytope covers through the training of several two-layer ReLU networks as per (5). It is crucial to note that all algorithms require a training dataset.
@@
  • The number of polytopes and faces. As detailed in Appendix C.2, Algorithm 3 generally yields multiple polytopes to correctly cover all data points present in the dataset. Additionally, it does not calculate the precise number of faces for each polytope. In contrast, Algorithm 2 generates a polytope-basis cover by compressing the given three-layer network. Therefore, if the three-layer network (6) is the sum of $J$ two-layer networks ($\mathcal{T}_{j}$), the algorithm is guaranteed to produce a polytope-basis cover consisting of no more than $J$ polytopes. The compressing algorithm and Lemma F.4 provide the exact number of faces for each polytope. Thirdly, Algorithm 4 does not impose any specific lower or upper bounds on the number of polytopes. Similar to Algorithm 2, it also furnishes the exact number of faces for each polytope.
@@
    For instance, we recall the example on the swiss roll dataset: Figures 2, 8, and 9 illustrate that the algorithms yield 16, 68, and 4 polytopes, respectively. Note also that the last algorithms suggest a feasible architecture of the dataset by $2\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}17\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}4\rightarrow 1$, while is sufficiently minimal.
@@
  • Theoretical guarantee. Proposition C.2 guarantees that Algorithm 2 and 3 must terminate in finite time. However, we do not provide theoretically guarantee for Algorithm 4. Even though, in all our experiments, it always terminates in finite time and produces a complete polytope-basis cover for the given dataset.
@@
  • Time consumption. Since Algorithm 3 does not require fine-tuning process, it consumes the shortest time among these algorithms. Algorithm 2 requires only fine-tuning process, so it takes a normal amount of time. However, Algorithm 4 tends to spend relatively more time due to its iterative process of training two-layer networks with fine-tuning until a complete polytope-basis cover is achieved. However, it’s essential to note that even for real-world datasets like CIFAR10, the practical execution time is still quite reasonable, typically taking only a few minutes (less than 15 minutes).
@@
  • Accuracy of the obtained cover. The polytope-basis cover generated by Algorithm 3 exhibits the same accuracy as the provided network, as confirmed by Proposition C.2. In the case of Algorithm 2, the fine-tuning process introduces the possibility of a different accuracy level compared to the original network. Consequently, the final accuracy cannot be determined beforehand. Regarding Algorithm 4, its accuracy is guaranteed to be 100% if it terminates within a finite time.
@@
Below, we provide additional experimental results of the proposed algorithms on several synthetic datasets, showcasing visual differences among these algorithms. We consider three synthetic datasets: XOR, two circles, and two moons datasets, depicted in Figure 10. The results of algorithms are shown in Figure 11, 12, 13.
@@
Figure 11: Visualization of Algorithm 2 on synthetic datasets. These polytope-basis covers are derived from trained three-layer ReLU networks.
Figure 12: Visualization of Algorithm 3 on synthetic datasets. These polytope-basis covers are derived from trained three-layer ReLU networks.
Figure 13: Visualization of Algorithm 4 on synthetic datasets. Empirically, this algorithm provides the smallest number of polytopes and their faces. The obtained polytope-basis covers can be applied to define the minimum architecture of neural networks.
It is easily checked that our proposed algorithms generate the polytope-basis cover of the given datasets. In each subfigure, the leftmost column represents the decision boundary and activation boundaries of the final network, and the other columns represent each polytope in the obtained polytope-basis cover. Note that the obtained polytope-basis covers exhibit the geometric characteristics of these datasets, and provide the minimal architecture of neural networks that can completely classify the given dataset.
For example, Figure 13 demonstrates that ’XOR’ and ‘two circles’ datasets have single polytope covers, and ‘two moons’ dataset can be covered by two polytopes. Therefore, the minimal network architecture of these dataset is given by

### v2

@@ C.2 A Polytope-Basis Covers Derived From a Two-Layer ReLU Network
Then, we quantize these functions to derive a polytope-basis cover. With the similar idea of Lebesgue integration, we can approximate the convex function $\mathcal{N}_{+}({\bm{x}})$ by a simple function. Here, the simple function means a linear combination of indicator functions, basically considered in mathematical field like Lebesgue theory (Rudin et al., 1976). Define $M:=\max_{{\bm{x}}\in{{\mathcal{D}}}}\mathcal{N}_{+}({\bm{x}})$. Then, for a given $\varepsilon>0$, we can approximate $\mathcal{N}_{+}({\bm{x}})$ by
@@
where $\mathcal{C}_{Q}$ is the collection of polytopes defined by $C_{l}:=\{{\bm{x}}~{}|~{}\mathcal{N}_{+}({\bm{x}})<M-l\varepsilon\}$ for $l=0,1,\cdots$. This decomposition can be understood as quantization of $\mathcal{N}_{+}({\bm{x}})$ by slices with height $\varepsilon$. Moreover, the above approximation would be accurate as $\varepsilon\rightarrow 0$. Similarly, $\mathcal{N}_{-}({\bm{x}})$ can be approximated in the similar manner. The main idea to obtain a polytope-basis cover $\mathcal{C}$ from $\mathcal{N}$ is selecting sufficiently many $\varepsilon$’s to quantize $\mathcal{N}_{+}$ and $\mathcal{N}_{-}$.
Empricially, we construct a polytope-basis cover from the values of $\mathcal{N}$. If $\mathcal{C}$ is not a polytope-basis cover yet, we select an incorrectly classified data point $\hat{{\bm{x}}}$ with the smallest confidence value, i.e., $\hat{\bm{x}}:=\operatorname*{arg\,min}_{{\bm{x}}_{i}\in{{\mathcal{D}}}}|\mathcal{N}({\bm{x}}_{i})|$. Then, we choose an intermediate value $c$ between $\mathcal{N}_{+}(\hat{{\bm{x}}})$ and $\mathcal{N}_{-}(\hat{{\bm{x}}})$ and add two polytopes $C_{+}:=\{{\bm{x}}~{}|~{}\mathcal{N}_{+}({\bm{x}})<c\}$ and $C_{-}:=\{{\bm{x}}~{}|~{}\mathcal{N}_{-}({\bm{x}})>-c\}$ to the polytope-basis cover $\mathcal{C}$. Then, the value
@@
        $\mathcal{C}\leftarrow\mathcal{C}_{P}\cup\mathcal{C}_{Q}$triangleright\;Now, $\hat{\bm{x}}$ is correctly covered by $\mathcal{C}$.
@@
The result of Algorithm 3 on a two-layer ReLU network with architecture $2\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}50\rightarrow 1$, which is trained on the Swiss roll dataset, is illustrated in Figure 9. Specifically, the algorithm generates a polytope-basis cover of the dataset consists of 68 polytopes ($\mathcal{C}_{P}$ and $\mathcal{C}_{Q}$ consists of 34 polytopes, respectively). 14 polytopes in $\mathcal{C}$ is illustrated in 2nd column to 8th column in Figure 9. Proposition C.2 guarantees that Algorithm 3 must terminate in finite time, and produces a polytope-basis cover of ${{\mathcal{D}}}$ which has the same accuracy with the given $\mathcal{N}$. Therefore, training a two-layer ReLU network to 100% accuracy on the dataset ${{\mathcal{D}}}$, Algorithm 3 allows to derive a polytope-basis cover of the given dataset.
There are some pros and cons in Algorithm 3. The advatages of the algorithm are 1. it can be applied to arbitrary two-layer ReLU networks, and 2. it does not modify the trained network. Therefore, it unveils the inherent polytope-basis cover and convex polytope structures in trained two-layer ReLU networks. However, two drawbacks in this algorithm are 1. generally it induces many polytopes in practice, and 2. the number of faces of each polytope is unknown. For detail comparison with other algorithms, see Section C.4.
C.3 An Efficient Algorithm to Find a Simple Polytope-Basis Cover
In the preceding subsections, we introduced two algorithms in Sections C.1 and C.2 that extract a polytope-basis cover from trained two-layer or three-layer ReLU networks. However, the results obtained from both algorithms, as illustrated in Figure 3 and 9, still exhibit too many polytopes on the training dataset. As demonstrated for the Swiss roll dataset in Figure 1, we have previously shown the existence of a polytope-basis cover comprising only four polytopes (refer to Figure 1 and 2).
@@ C.3 An Efficient Algorithm to Find a Simple Polytope-Basis Cover
Before we demonstrate the algorithm, we provide modified network and loss functions. Specifically, we consider the following two types of two-layer ReLU networks. Here, $\lambda_{bias}>0$ is a hyperparameter that enhances the convergences as introduced in (5).
@@
If $\lambda_{bias}$ is large, then $\mathcal{T}_{+}$ has a large output value at initialization, and gradient descent optimization is heavily affected by data with 00 labels. A similar situation happens for $\mathcal{T}_{-}$, and it helps to find a single polytope that contains whole class data. Practically, it is enough to use $\lambda_{bias}=5$ to find such polytopes.
@@
where $\bm{\lambda}=(\lambda_{0},\lambda_{1})$ is the hyperparameter proposed to reinforce to cover a whole data class with specific label. For instance, by using a large value of $\lambda_{1}$, $\mathcal{T}$ can be trained to cover whole data points that have label $y_{i}=1$. After successfully configured the first polytope $C_{1}$, we train the second network $\mathcal{T}_{2}$ to distinguish data points of another data class, $y_{i}=0$, in the obtained polytope $C_{1}$. Repeating this process alternatively, Algorithm 4 generally provides a polytope-basis cover with a small number of polytopes.
Now, we provide a detailed illustration of the algorithm’s functionality with the example displayed in Figure 10. Here we use $\lambda_{bias}=5$, and $\bm{\lambda}=(1,10)$. First, the algorithm trains $\mathcal{T}_{1}$ on the entire dataset ${{\mathcal{D}}}$ using the loss function in (12). After fine-tuning through Algorithm 2, we obtain the first polytope $C_{1}$ displayed in Figure 10(b). Note that all orange data points $({{\mathcal{D}}}_{1})$ are contained in $C_{1}$. Next, the second network $\mathcal{T}_{2}$ is trained on $({{\mathcal{D}}}_{0}\cap C_{1})\cup{{\mathcal{D}}}_{1}$ using (12) with $\bm{\lambda}=(1,10)$. $\mathcal{T}_{2}$ is aimed to cover all blue data points $({{\mathcal{D}}}_{0})$ within $C_{1}$. After training and fine-tuning, we obtain the second polytope $C_{2}$ displayed in Figure 10(c). Similarly, we can find the third polytope $C_{3}$ by training another network $\mathcal{T}_{3}$ on ${{\mathcal{D}}}_{0}\cup({{\mathcal{D}}}_{1}\cap C_{2})$, displayed in Figure 10(d) and so on. In the example in Figure 10, totally four polytopes are obtained, and visualized through (b) to (e). Lastly, by Theorem 3.4, we can construct a three-layer ReLU network with architecture $2\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}17\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}4\rightarrow 1$ that can completely classify this swiss roll dataset, based on the obtained polytope-basis cover. The decision boundary of the constructed three-layer ReLU network is illustrated in Figure 10(f).
Algorithm 4 An efficient algorithm for finding a polytope-basis cover
@@
The key advantage of this algorithm lies in its ability to generate a small number of convex polytopes, in contrast to other algorithms we proposed. As demonstrated in Figure 10, it suggests $2\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}17\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}4\rightarrow 1$ as a feasible architecture on the given Swiss roll dataset, which appears to be close to optimal. However, due to the iterative nature of training multiple two-layer networks until achieving a complete polytope-basis cover, it typically requires a longer computation time. A detailed comparison with other algorithms is provided in the subsequent section.
Figure 10: The result of Algorithm 4. For the given dataset (a), the algorithm determines the first polytope $C_{1}$ which contains the whole orange class, as shown in (b). In the next step, it obtains the second polytope $C_{2}$ which contains the whole blue class inside $C_{1}$. Other polytopes are similarly derived and illustrated in (c) to (e). Totally, the algorithm produces a polytope-basis cover consisting of four polytopes. Theorem 3.4 shows that $2\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}17\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}4\rightarrow 1$ is a feasible architecture on this dataset, and the decision boundary of the induced network is drawn in (f).
C.4 Comparison of the Proposed Algorithms
Finally, we compare the proposed algorithms (Algorithm 1, 2, 3, and 4). First, we provide theoretical results for the proposed algorithms, where their proofs involve demonstrating how a polytope-basis cover can be explicitly constructed from the results of the algorithms. The detailed proof is available in Appendix E.6.
@@ C.4 Comparison of the Proposed Algorithms
    Algorithm 2 must terminate in finite time, and it produces a polytope-basis cover of the training dataset ${{\mathcal{D}}}$ that has the same accuracy with the compressed network $\mathcal{N}$.
@@
    Algorithm 3 must terminate in finite time, and it produces a polytope-basis cover of the training dataset ${{\mathcal{D}}}$ that has the same accuracy with the given network $\mathcal{N}$.
@@
# of obtained polytopes at most $J$ generally large generally small
@@
Accuracy of the polytope cover unknown same with the given network $>acc_{th}$ (if it terminates)
Table 3: Comparison of proposed algorithms. All these algorithms generate a polytope-basis cover of the given training dataset, but each of them has its own pros and cons.
@@
  • Input network. Algorithm 2 operates on a pretrained three-layer network outlined in (6). Algorithm 3 necessitates a fully-trained two-layer ReLU network specified by (1). However, Algorithm 4 does not necessitate any pre-existing networks as inputs but generates polytope covers through the training of several two-layer ReLU networks as per (5).
@@
  • The number of polytopes and faces. As detailed in Appendix C.2, Algorithm 3 generally yields multiple polytopes to correctly cover all data points in the dataset. Additionally, it does not calculate the precise number of faces for each polytope. In contrast, Algorithm 2 generates a polytope-basis cover by compressing the given three-layer network. Therefore, if the three-layer network $\mathcal{N}$ defined in (6) is the sum of $J$ two-layer networks ($\mathcal{T}_{j}$), the algorithm is guaranteed to produce a polytope-basis cover consisting of no more than $J$ polytopes. The compressing algorithm and Lemma F.4 provide the exact number of faces for each polytope. Thirdly, Algorithm 4 does not impose any specific lower or upper bounds on the number of polytopes. Similar to Algorithm 2, it also furnishes the exact number of faces for each polytope.
@@
    For instance, we recall the example on the swiss roll dataset: Figures 3, 9, and 10 illustrate that the algorithms yield 16, 68, and four polytopes, respectively. Also note that the last algorithms suggest a feasible architecture of the dataset by $2\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}17\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}4\rightarrow 1$, which looks sufficiently minimal.
@@
  • Theoretical guarantee. Proposition C.2 guarantees that Algorithm 2 and 3 must terminate in finite time. However, there is no theoretically guarantee for Algorithm 4. Even though, in all our experiments on synthetic and real-world datasets, it always terminates in finite time and produces a complete polytope-basis cover for the given dataset.
@@
  • Time consumption. Since Algorithm 3 does not require fine-tuning process, it consumes the shortest time among these algorithms. Algorithm 2 requires only fine-tuning process, so it takes a normal amount of time. However, Algorithm 4 tends to spend relatively longer time due to its iterative process of training two-layer networks with fine-tuning until it achieves a complete polytope-basis cover. However, it is essential to note that even for real-world datasets like CIFAR10, the practical execution time is still quite reasonable, typically taking only a few minutes.
@@
  • Accuracy of the obtained cover. The polytope-basis cover generated by Algorithm 3 preserves the accuracy of the given network (Proposition C.2). In the case of Algorithm 2, the fine-tuning process introduces the possibility of a different accuracy level compared to the original network. Consequently, the final accuracy cannot be determined beforehand. Regarding Algorithm 4, its accuracy is guaranteed to be greater than the given $acc_{th}$ if it terminates within a finite time.
@@
Below, we provide additional experimental results of the proposed algorithms on several synthetic datasets, showcasing visual differences among these algorithms. We consider three synthetic datasets: XOR, two circles, and two moons datasets, depicted in Figure 11. The results of algorithms are shown in Figure 12, 13, and 14.
@@
Figure 12: Visualization of Algorithm 2 on the synthetic datasets. These polytope-basis covers are derived from trained three-layer ReLU networks (6) with the architecture $2\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}80\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}8\rightarrow 1$ (i.e., a combination of eight two-layer networks $\mathcal{T}_{j}$ with $100pt=$ 10 neurons)
Figure 13: Visualization of Algorithm 3 on the synthetic datasets. These polytope-basis covers are derived from trained two-layer ReLU networks with the architecture $2\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}20\rightarrow 1$.
Figure 14: Visualization of Algorithm 4 on the synthetic datasets. Empirically, this algorithm provides the smallest number of polytopes and their faces. The obtained polytope-basis covers can be applied to conclude the feasible architecture of neural networks (Remark C.3).
It is easily checked that our proposed algorithms indeed generate polytope-basis covers of the given datasets. In each subfigure, the leftmost column represents the decision boundary and activation boundaries of the obtained networks, and the other columns represent each polytope in the obtained polytope-basis cover. The obtained polytope-basis covers exhibit the geometric characteristics of datasets, and provide feasible architectures of neural networks.
Remark C.3.
Figure 14 demonstrates that ’XOR’ and ‘two circles’ datasets have single polytope covers, and ‘two moons’ dataset can be covered by two polytopes. From the obtained polytope-basis covers, the feasible architectures of these datasets are given by